An insulated water heater operating at steady state has two inlets and one 

outlet. At inlet 1, liquid water enters at 7 bar, 42oC, and a mass flow rate of 70 

kg/s. At inlet 2, wet vapor at 7 bar with a quality of 98% is injected. At outlet, 

saturated liquid at 7 bar exits the water heater. Ignoring heat transfer with the 

surroundings and neglecting kinetic and potential energy effects, determine 

the mass flow rate, in kg/s, at inlet 2


1
Expert's answer
2021-07-07T11:39:44-0400

Answer:-

At inlet 1 ,

"P_1= 7 \ bar \\
T_1 =42^o C \\
m_1=70 kg/s"

At inlet 2

"P_2=7 \ bar\\
x_2=0.98\\"

at exit 3 , "p_3=7 \ bar"



at T1 = 42o C from saturated stream tables ,

hf1=167.57 + (188.45-167.57) "[\frac{42-40}{45-40}]=175.9kJ/kg"

"V_{f1}=1.0078\times 10^{-3}+(1.0099\times10^{-3}-1.0078\times10^{-3})\times [\frac{42-40}{45-40}]\\=1.0086\times10^{-3}m^3/kg"

"P_{Sat1}=0.07384+(0.09593-0.07384)\times [\frac{42-40}{45-40}]
\\=0.08268 bar"


At "P_2=7" bar from saturated steam tables,

"h_{f2}=697.22 kJ/kg\\
h_{fs2}=2066.3 KJ/Kg\\
h_2=h_{f2}+X_2h_{fs2}"

"h_2=697.22+0.98\times2066.3\\
h_22722.2 KJ/Kg"


at "p_3 =7 \ bar" from saturated steam tables ,

"h_3=h_{f3}"

"h_3=697.22KJ\/kg" ("\therefore" saturated liquid state)


By mass rate balance under steady state

"m_1+m_2+m_3=0.....(1)"

By energy rate balance equation

"0=Q_{cv}-W_{cv}+m_1(h_1+\frac{v_1^2}{2}+gz_1)+m_2(h_2+\frac{v_2^2}{2}+gz_2)-m_3(h_3+\frac{v_3^2}{2}+gz_3)"

"\therefore" work done , heat transfer , kinetic and

potential energies are negligible

"m_1h_1+m_2h_2+m_3h_3=0...........(2)"


using (1) in (2)

"m_1h_1+m_2h_2-(m_1+m_2)h_3=0\\"

on putting the values

"m_2=70\times\frac{(697.22-176.6)}{(2722.2-697.22)}"

"m_2=18 \ kg/s"

Mass flow rate 2 is "m_2=18 \ kg/s"